# Bidirectional Task-Motion Planning Based on Hierarchical Reinforcement Learning for Strategic Confrontation

### v2

@@ III-C Embedded Trajectory Prediction Model
When the agent performs a searching or escaping task, it invokes a trajectory prediction model to compute the subgoal. In this paper, we assume that a global planner stores the historical trajectories of all enemies $p_{e,k}\left(t\right),t=0,\dots,T_{n}$, where $T_{n}$ represents the current time step. The planner then makes a preliminary trajectory prediction using a second-order dynamic model
$\begin{array}[]{c}\boldsymbol{p}_{e,k}\left(T_{n}+1\right)=\boldsymbol{p}_{e,k}\left(T_{n}\right)+\delta t\boldsymbol{v}_{e,k}\left(T_{n}\right)+\frac{1}{2}\delta t^{2}\boldsymbol{a}_{e,k}\left(T_{n}\right),\end{array}$ (14)
where $\boldsymbol{a}_{e,k}\left(T_{n}\right)=\left(\boldsymbol{v}_{e,k}\left(T_{n}\right)-\boldsymbol{v}_{e,k}\left(T_{n}-1\right)\right)/\delta t$. To avoid collisions with obstacles in the confrontation scenario, the planner performs collision detection for each trajectory and further refines the trajectory through the potential field method. We consider an obstacle circle $c\in\mathbb{N}_{k}$ if it satisfies $\lVert\boldsymbol{p}_{e,k}\left(T_{n}+1\right)-\boldsymbol{p}_{o,c}\rVert<\rho_{2}+\rho_{3}$, where $\mathbb{N}_{k}$ denotes the set of obstacles that collide with the enemy $k$. The predicted position is then adjusted as follows:
@@
where $d_{3}=\rho_{2}+\rho_{3}-\lVert\boldsymbol{p}_{e,k}\left(T_{n}+1\right)-\boldsymbol{p}_{o,c}\rVert$. After iterating until time step $T_{n}+T_{p}$, the planner clusters all predicted trajectory points using a hierarchical clustering algorithm [23] and determines the center of the $b_{1}$-th cluster, which represents one of the most likely locations of enemies:
@@
where $\boldsymbol{p}_{n}$ is the position of each point in the cluster, and $N_{p}$ is the number of points in the cluster. The symbol $T_{p}$ is a constant that indicates the length of the prediction time step. If there are no enemies within the perception range, and no allies in chasing or escaping within the communication range, the $i$-th agent will switch to a searching task, selecting the nearest center as its subgoal
@@
Otherwise, if an enemy is within the perception range of the $i$-th agent and satisfies $Adv\left(\boldsymbol{u}_{i},\boldsymbol{u}_{e,k}\right)<0$, the agent will perform an escaping task. To compute the least likely locations of enemies, the planner randomly generates several candidate points that satisfy the following conditions:
@@
The subgoal is then transmitted from the global planner to the $i$-th agent.
Figure 3: Interaction mechanism of two-layer networks.
@@ III-D Training Methods for Two Layers
Combining the goal-conditioned MDP framework, we adopt double DQN for training in the upper layer. In this approach, each agent estimates the optimal state-action value function $Q_{i}^{*}:\mathbb{S}\times\mathbb{A}\times\mathbb{G}\rightarrow\mathbb{R}$ through a parameterized neural network $Q_{\phi_{u,i}}\left(\boldsymbol{s}_{i}\left(t\right),\boldsymbol{a}_{i}\left(t\right),\boldsymbol{g}\right)$. The subscript $\phi_{u}=\left(\phi_{u,1},\dots,\phi_{u,N}\right)$ represents the weight parameters of the value-function network in the upper layer. For $\gamma\approx 1$, $Q^{*}$ estimates the discounted returns of the optimal strategy over an infinite horizon. The method approximates $Q^{*}$ using the network $Q_{\phi_{u,i}}$, and the loss function is formulated as
@@
where $E$ denotes the mathematical expectation. The $Q$-target $y_{i}$ is defined as
@@
where the subscript $\phi_{l}=\left(\phi_{l,1},\dots,\phi_{l,N}\right)$ represents the weight parameters of the critic network in the lower layer. The $Q$-target is revised as
@@
Moreover, we propose a cross-training method to address the strong correlation between the two layers, as shown in Algorithm 1. It updates the upper and lower networks simultaneously, enabling both layers to adjust their policies in real time. This method prevents training instability due to the interactions between the two layers.
Algorithm 1 Cross-Training Method
0:  the number of training steps $T$; the number of cross-training episodes $E_{c}$; the length of each subtask $h$;
0:  the training datasets of the training processes; the parameters of the networks in the upper layer $\phi_{u}$ and lower layer $\phi_{l},\omega_{l}$;
[…]
@@ IV Experiments
In this section, extensive experiments are conducted to evaluate the performance of our HRL method in solving the strategic confrontation problem at different sizes. In comparative experiments, we adopt various baselines, including the expert system, heuristic algorithm, and single-layer DRL. We assess the impact of the interaction mechanism through ablation studies. Moreover, we test the model trained on a small scale to solve larger-scale problems, investigating the generalization of our method. Finally, we deploy the cross-trained model from simulations to the real-robot system.
@@ IV-A Setting Up
Before analyzing the comparative results, we outline the detailed settings used in the simulations. Following the conventions in [8, 22], we consider an equal number of red and blue agents, randomly distributed in a 30m×$\times$20m scenario with static obstacles. In our experiment, three scenarios are considered: five agents versus five agents (denoted V5), seven agents versus seven agents (denoted V7), and nine agents versus nine agents (denoted V9). The attack radius $\rho_{1}$, angle $\theta_{1}$, and probability $\epsilon$ are set to 1m, $\pi/$2, and 0.5, respectively. The avoidance radius $\rho_{2}$ and the obstacle radius $\rho_{3}$ are set to 0.3m. The observation length $d_{1}$ and width $d_{2}$ are set to 5m. The maximum velocity magnitude $\lVert\boldsymbol{v}\rVert_{\text{max}}$ is 2m/s, and the step length $\delta t$ is 0.1s.
For algorithm training, we cross-train the two layers for $E_{c}=$ 60 episodes and $T=$ 300 steps in 100 randomly generated instances. The discount factor $\gamma$ and the learning rate $\alpha$ are set to $0.99$ and $0.001$, respectively. To plot the experimental curves, we use solid curves to represent the mean values across all instances, with shaded regions indicating the standard deviation. In our experiments, we apply the HRL method and baselines to the blue agents, while the red agents use traditional TAMP methods, including the expert system and heuristic algorithm.
@@ IV-B Comparative Analysis
In cross-training, we employ the HRL method to enhance the performance of both layers. We compare this method with baselines, including the expert system, heuristic approach, and single-layer DRL. Additionally, to investigate the influence of interactions between the two layers, we conduct an ablation study by removing the feedback reward in HRL. The baselines are briefly introduced as follows.
@@
    Expert system: Based on the confrontation rules constructed in [8], this algorithm uses a hierarchical decision-making framework to match optimal task allocation and path planning.
@@
    Heuristic algorithm: The algorithm in [11] designs an evolutionary-based attack strategy for swarms, evaluating the potential benefits or threats of the next perceptible attack positions of an agent. It also incorporates a collision-avoidance method for feasible path planning.
@@
    Single-layer DRL: The algorithm in [17] treats confrontation as an MDP process and employs single-layer DRL to output the agents’ strategies end-to-end.
@@
enemies Re. Ti.(×$\times 10^{-2}$s) W.R.(%) Re. Ti.(×$\times 10^{-2}$s) W.R.(%) Re. Ti.(×$\times 10^{-2}$s) W.R.(%)
@@
To ensure a fair comparison between our method and the baselines, we evaluate all methods under the same experimental conditions, using identical input data and performance metrics. The HRL learning curves and the baselines for different swarm sizes are shown in Fig. 4, where "VX-E" and "VX-H" represent the use of the expert system and heuristic algorithm as enemies’ strategies in the VX scenario, respectively. Both the expert system and heuristic algorithm can develop effective strategies for agents across different instances. Between these two conventional TAMP algorithms, the heuristic algorithm outperforms the expert system in large-scale confrontations.
@@
Initially, learning-based algorithms underperform due to the randomness of their starting strategies. With continuous exploration and training, their performance gradually surpasses that of the non-learning algorithms. Due to the hybrid decision-making process in the confrontation, the performance of single-layer DRL deteriorates over time during training. In contrast, HRL consistently finds better strategies than non-learning algorithms across various swarm sizes. Furthermore, when the feedback reward is omitted, the performance curve converges to a suboptimal value due to the lack of effective interaction between the two layers.
@@
In addition, we deploy the policy networks trained using learning-based methods across another 100 instances and compare their performance with conventional TAMP algorithms. To provide a more comprehensive assessment, we introduce two additional evaluation metrics: decision time and confrontation win rate. Decision time refers to the average time that all agents take to allocate tasks and plan the path during each time step. Confrontation win rate measures the ratio of successful outcomes to the total number of instances. The average experimental results across 100 instances for different swarm sizes are summarized in Table I, which presents the episode returns (Re.), decision time (Ti.), and confrontation win rate (W.R.) for all methods.
@@
Since the decision time for all methods is shorter than the time step $\delta t$, these methods are capable of real-time decision-making at each step. We observe a clear relationship between the win rate and episode returns, which are calculated using the reward function in this study. Agents train their strategies to maximize returns, which in turn leads to a higher win rate.
@@
Between the two non-learning algorithms, the heuristic algorithm has a longer decision time, though its returns and win rate improve as the problem size increases. Due to its end-to-end nature, DRL exhibits the shortest decision time, but its win rate is significantly lower than that of HRL. The results demonstrate that our method outperforms the baselines in terms of episode returns, decision time, and confrontation win rate, especially in large-scale instances. This is attributed to the effectiveness of our approach in managing interactions between the upper and lower layers by decoupling confrontation into allocation and planning.
Figure 5: Generalization to larger-scale instances.
@@
IV-C Generalization to Larger-Size Instances
@@
In large-scale confrontation scenarios, retraining strategies can lead to significant consumption of computational resources. Therefore, our goal is for the trained models to exhibit strong generalization performance. In our method, the decentralized policies of agents can be applied directly to additional agents without the need for retraining. To evaluate the generalization capability of our approach, we apply the two-layer networks trained on the V9 scenario to instances with more agents (ranging from 10 to 20). As the scenario size increases, agents face greater challenges and experience lower win rates. However, our method continues to maintain favorable solution quality compared to the baselines.
@@
The generalization results are shown in Fig. 5, which presents the win rate across different instance scales. Since traditional TAMP algorithms lack generalization capabilities, we reiterate them for larger-scale scenarios. As shown, our method outperforms expert system, heuristic algorithm, and single-layer DRL in most instances. Our approach demonstrates high scalability in larger scenarios compared to the baselines. In particular, the model trained on the V9 scenario can generalize to V20, achieving win rates consistent with those achieved by the heuristic algorithm reiterated on V20. These results verify that policies trained on small-scale instances can effectively make decisions in scenarios of varying scales. Thus, our method exhibits favorable generalization capability for the strategic confrontation problem.
IV-D Deployment in Real-Robot System
@@
Before presenting the real-robot experiments, we evaluate the performance of our method through simulations that account for communication and sensor uncertainties, including message loss between neighbors and sensor noise error. The results are shown in Fig. 6, where the horizontal axis represents the message loss rate, the vertical axis indicates the win rate, and the legend represents different levels of Gaussian noise error. Although the success rate decreases as the uncertainty increases, it remains above 73% when the loss rate is below 10% and the noise error is within ±5%. These results demonstrate that our method is suitable for real-world applications.
@@
We then conduct experiments using an actual robotic system, as depicted in Fig. 7. A V3 confrontation is considered in a static-obstacle scenario, with experimental settings consistent with those defined in Section II-A. By performing cross-training in simulations, we obtain two-layer networks for the agents in this scenario. During the experiment, we control the agents via a ground control system, with motion capture provided by Optitrack. Our algorithm receives data from Optitrack that includes the positions and velocities of all agents. The algorithm subsequently calculates observations from perception sensors and generates position commands based on the two-layer networks, which are sent to the ground control system. Using these networks, the agents successfully defeat all enemies during the experiments while avoiding collisions. After sufficient training with our method, the agents exhibit coordinated confrontation behaviors, such as prioritizing support for teammates in critical situations (Fig. 7(a)), using static obstacles for cover (Fig. 7(b)), and luring opponents into teammates’ encirclement (Fig. 7(c)).
@@
Figure 7: The decision-making process for V3 scenario in the real-robot system.

### v3

@@ III-C Embedded Trajectory Prediction Model
When the agent performs a searching or escaping task, it invokes a trajectory prediction model to compute the subgoal. In this paper, we assume that a global planner stores the historical trajectories of all enemies $p_{e,k}\left(t\right),t=0,\dots,T_{n}$, where $T_{n}$ represents the current time step. The planner then makes a preliminary trajectory prediction using a second–order dynamic model
$\begin{array}[]{c}\boldsymbol{p}_{e,k}\left(T_{n}+1\right)=\boldsymbol{p}_{e,k}\left(T_{n}\right)+\delta t\boldsymbol{v}_{e,k}\left(T_{n}\right)+\frac{1}{2}\delta t^{2}\boldsymbol{\dot{v}}_{e,k}\left(T_{n}\right),\end{array}$ (13)
where $\boldsymbol{\dot{v}}_{e,k}\left(T_{n}\right)=\left(\boldsymbol{v}_{e,k}\left(T_{n}\right)-\boldsymbol{v}_{e,k}\left(T_{n}-1\right)\right)/\delta t$. To avoid collisions with obstacles in the confrontation scenario, the planner performs collision detection for each trajectory and further refines the trajectory through the potential field method. We consider an obstacle circle $c\in\mathbb{N}_{k}$ if it satisfies $\lVert\boldsymbol{p}_{e,k}\left(T_{n}+1\right)-\boldsymbol{p}_{o,c}\rVert<\rho_{2}+\rho_{3}$, where $\mathbb{N}_{k}$ denotes the set of obstacles that collide with the enemy $k$. The predicted position is then adjusted as follows:
@@
where $d_{3}=\rho_{2}+\rho_{3}-\lVert\boldsymbol{p}_{e,k}\left(T_{n}+1\right)-\boldsymbol{p}_{o,c}\rVert$. After iterating until time step $T_{n}+T_{p}$, the planner employs a hierarchical clustering algorithm [23] to group all predicted trajectory points. To generate trajectories for clustering, we repeatedly roll out the prediction model from each enemy’s most recent position. The resulting positions are then used for hierarchical clustering. The center of the $b_{1}$–th cluster, representing one of the most likely locations of enemies, is given by
@@
where $\boldsymbol{p}_{n}$ is the position of each point in the cluster, and $N_{p}$ is the number of points in the cluster. The symbol $T_{p}$ is a constant that indicates the length of the prediction time step. If there are no enemies within the perception range, and no allies in chasing or escaping within the communication range, the ii–th agent will switch to a searching task, selecting the nearest center as its subgoal
@@
Otherwise, if an enemy is within the perception range of the ii–th agent and satisfies $Adv\left(\boldsymbol{u}_{i},\boldsymbol{u}_{e,k}\right)<0$, the agent will perform an escaping task. To compute the least likely locations of enemies, the planner randomly generates several candidate points that satisfy the following conditions:
@@
The subgoal is then transmitted from the global planner to the ii–th agent.
Figure 3: Interaction mechanism of two–layer networks.
@@ III-D Training Methods for Two Layers
Combining the goal–conditioned MDP framework, we adopt double DQN for training in the upper layer. In this approach, each agent estimates the optimal state–action value function $Q_{i}^{*}:\mathbb{S}\times\mathbb{A}\times\mathbb{G}\rightarrow\mathbb{R}$ through a parameterized neural network $Q_{\phi_{u,i}}\left(\boldsymbol{s}_{i}\left(t\right),\boldsymbol{a}_{i}\left(t\right),\boldsymbol{g}\right)$. The subscript $\phi_{u}=\left(\phi_{u,1},\dots,\phi_{u,N}\right)$ represents the weight parameters of the value–function network in the upper layer. For $\gamma\approx 1$, $Q^{*}$ estimates the discounted returns of the optimal strategy over an infinite horizon. The method approximates $Q^{*}$ using the network $Q_{\phi_{u,i}}$, and the loss function is formulated as
@@
where $E$ denotes the mathematical expectation. The QQ–target $y_{i}$ is defined as
@@
where the subscript $\phi_{l}=\left(\phi_{l,1},\dots,\phi_{l,N}\right)$ represents the weight parameters of the critic network in the lower layer. The QQ–target is revised as
@@
To ensure stable policy convergence under the strong coupling between the two layers, we introduce a cross–training method. The training alternates between updating the lower–layer policies every h steps and subsequently refining the upper–layer policy using the accumulated feedback. Details are outlined in Algorithm 1. This method updates the upper and lower networks simultaneously, enabling both layers to adjust their policies in real time.
Algorithm 1 Cross–Training Method
0: the number of training steps $T$; the number of cross–training episodes $E_{c}$; the length of each subtask $h$;
0: the parameters of the networks in the upper layer $\phi_{u}$ and lower layer $\phi_{l},\omega_{l}$;
1:for all $e\ =\ 1\rightarrow E_{c}$ do
@@ IV Experiments
In this section, extensive experiments are conducted to evaluate the performance of our HRL method in solving the strategic confrontation problem at different sizes. In comparative experiments, we adopt various baselines, including the expert system, heuristic algorithm, and single–layer DRL. We assess the impact of the interaction mechanism through ablation studies. Moreover, we test the model trained on a small scale to solve larger–scale problems, investigating the generalization of our method. Finally, we deploy the cross–trained model from simulations to the real–robot system.
@@ IV-A Setting Up
[…]
For algorithm training, we cross–train the two layers for $E_{c}=$ 60 episodes and $T=$ 300 steps in 100 randomly generated instances. The networks in both layers are implemented as three–layer MLPs with hidden units [128, 64, 32]. All networks use ReLU activations and Adam optimizer. The replay buffer size is 105, and mini–batch size is 64. The discount factor $\gamma$ and the learning rate $\alpha$ are set to $0.99$ and $0.001$, respectively. To plot the experimental curves, we use solid curves to represent the mean values across all instances, with shaded regions indicating the standard deviation. In our experiments, we apply the HRL method and baselines to the blue agents, while the red agents use traditional TAMP methods, including the expert system [8] and heuristic algorithm [11].
@@ IV-B Comparative Analysis
In cross–training, we employ the HRL method to enhance the performance of both layers. We compare this method with baselines, including the expert system, heuristic approach, and single–layer DRL. Additionally, to investigate the influence of interactions between the two layers, we conduct an ablation study by removing the feedback reward in HRL. The baselines are briefly introduced as follows.
@@
    Expert system: Based on the confrontation rules constructed in [8], this algorithm uses a hierarchical decision–making framework to match optimal task allocation and path planning.
@@
    Heuristic algorithm: The algorithm in [11] designs an evolutionary–based attack strategy for swarms, evaluating the potential benefits or threats of the next perceptible attack positions of an agent. It also incorporates a collision–avoidance method for feasible path planning.
@@
    Single–layer DRL: The algorithm in [17] treats confrontation as an MDP process and employs single–layer DRL to output the agents’ strategies end–to–end.
@@
enemies Re. Ti.($\times 10^{-2}$ s) W.R.(%) Re. Ti.($\times 10^{-2}$ s) W.R.(%) Re. Ti.($\times 10^{-2}$ s) W.R.(%)
@@
To ensure a fair comparison between our method and the baselines, we evaluate all methods under the same experimental conditions, using identical input data and performance metrics. The HRL learning curves and the baselines for different swarm sizes are shown in Fig. 4, where "VX–E" and "VX–H" represent the use of the expert system and heuristic algorithm as enemies’ strategies in the VX scenario, respectively. Both the expert system and heuristic algorithm can develop effective strategies for agents across different instances. Between these two conventional TAMP algorithms, the heuristic algorithm outperforms the expert system in large–scale confrontations.
@@
Initially, learning–based algorithms underperform due to the randomness of their starting strategies. With continuous exploration and training, their performance gradually surpasses that of the non–learning algorithms. Due to the hybrid decision–making process in the confrontation, the performance of single–layer DRL deteriorates over time during training. In contrast, HRL consistently finds better strategies than non–learning algorithms across various swarm sizes. Furthermore, when the feedback reward is omitted, the performance curve converges to a suboptimal value due to the lack of effective interaction between the two layers.
@@
In addition, we deploy the policy networks trained using learning–based methods across another 100 instances and compare their performance with conventional TAMP algorithms. To provide a more comprehensive assessment, we introduce two additional evaluation metrics: decision time and confrontation win rate. Decision time refers to the average time that all agents take to allocate tasks and plan the path during each time step. Confrontation win rate measures the ratio of successful outcomes to the total number of instances. The average experimental results across 100 instances for different swarm sizes are summarized in Table I, which presents the episode returns (Re.), decision time (Ti.), and confrontation win rate (W.R.) for all methods.
@@
Notably, all methods support real–time decision–making at each step, with decision time below the time interval $\delta t$. We observe a clear relationship between the win rate and episode returns, which are calculated using the reward function in this study. Agents train their strategies to maximize returns, which in turn leads to a higher win rate.
@@
Between the two non–learning algorithms, the heuristic algorithm has a longer decision time, though its returns and win rate improve as the problem size increases. Due to its end–to–end nature, DRL exhibits the shortest decision time, but its win rate is significantly lower than that of HRL. The results demonstrate that our method outperforms the baselines in terms of episode returns, decision time, and confrontation win rate, especially in large–scale instances. This is attributed to the effectiveness of our approach in managing interactions between the upper and lower layers by decoupling confrontation into allocation and planning.
Figure 5: Generalization to larger–scale instances.
@@
IV-C Generalization to Larger–Size Instances
@@
In large–scale confrontation scenarios, retraining strategies can lead to significant consumption of computational resources. Therefore, our goal is for the trained models to exhibit strong generalization performance. In our method, the decentralized policies of agents can be applied directly to additional agents without the need for retraining. To assess the generalization capability of our approach, we apply a set of neural network parameters, trained on the V9 scenario, to instances involving a higher number of agents (ranging from 10 to 20). As the scenario size increases, agents face greater challenges and experience lower win rates. However, our method continues to maintain favorable solution quality compared to the baselines.
@@
The generalization results are shown in Fig. 5, which presents the win rate across different instance scales. Due to the limited generalization capabilities of traditional TAMP algorithms, they require repeated adjustments for larger–scale scenarios. As shown, our method outperforms expert system, heuristic algorithm, and single–layer DRL in most instances. Our approach demonstrates high scalability in larger scenarios compared to the baselines. In particular, the model trained on the V9 scenario can generalize to V20, achieving win rates consistent with those achieved by the heuristic algorithm reiterated on V20. These results verify that policies trained on small–scale instances can effectively make decisions in scenarios of varying scales. Thus, our method exhibits favorable generalization capability for the strategic confrontation problem.
IV-D Deployment in Real–Robot System
@@
Before presenting the real–robot experiments, we evaluate the performance of our method through simulations that account for communication and sensor uncertainties, including message loss between neighbors and sensor noise error. The results are shown in Fig. 6, where the horizontal axis represents the message loss rate, the vertical axis indicates the win rate, and the legend represents different levels of Gaussian noise error. Although the success rate decreases as the uncertainty increases, it remains above 73% when the loss rate is below 10% and the noise error is within ±5%. These results demonstrate that our method is suitable for real–world applications.
@@
We then conduct experiments using an actual robotic system, as depicted in Fig. 7. A V3 confrontation is considered in a static–obstacle scenario, with experimental settings consistent with those defined in Section II-A. By performing cross–training in simulations, we obtain two–layer networks for the agents in this scenario. During the experiment, we control the agents via a ground control system, with motion capture provided by Optitrack. Our algorithm receives data from Optitrack that includes the positions and velocities of all agents. The algorithm subsequently calculates observations from perception sensors and generates position commands based on the two–layer networks, which are sent to the ground control system. Using these networks, the agents successfully defeat all enemies during the experiments while avoiding collisions. After sufficient training with our method, the agents exhibit coordinated confrontation behaviors, such as prioritizing support for teammates in critical situations (Fig. 7(a)), using static obstacles for cover (Fig. 7(b)), and luring opponents into teammates’ encirclement (Fig. 7(c)).
@@
Figure 7: The decision–making process for V3 scenario in the real–robot system.